# GW/DT invariants and 5D BPS indices for strips from topological recursion

### v1

@@ 3.1. Four punctured spheres
Depending on whether $|Q|<1$ or $|Q|>1$, one can write the sum in terms of $\mathrm{Li}_{3-2g}(Q)$ or $\mathrm{Li}_{3-2g}(Q^{-1})$ which are equal. However, as expansions in $Q$, only one of them is convergent, given the above choice. This was also observed in [Sze08], giving two distinct chambers in complex structure moduli space. In this case, we infer
@@
This matches with the results of [OSY11, Sze08, BLR21].
@@
$Z_{\mathrm{DT}}=M(q)^{-2}\prod_{n=1}^{\infty}(1-\beta q^{n})^{-n}$ (3.6)
equation (4.36) in [BLR21].
@@ 3.2. Examples of five punctured spheres
Next, one could consider the case of gluing two copies of $\mathcal{O}(0,-2)$, corresponding to $s=2$ which corresponds to resolution of $\mathbb{C}^{3}/\mathbb{Z}_{3}$. In this case, as in [PS19], we consider $\beta_{1}=Q_{1},\beta_{2}=Q_{1}Q_{2}$. Looking at the genus $g$ free energy (2.15), one can check
@@
which matches the computations of [CMPS17, MP25, Mor12, MN15].
@@ 3.3. Examples of six punctured spheres
We take the example of [PS19] for the six-punctured sphere with $r=2,s=1$, with $\alpha_{1}=Q_{1},\alpha_{2}=Q_{1}Q_{2}Q_{3}$ and $\beta_{1}=Q_{1}Q_{2}$. The 5D indices can then be read off from (2.15) as
@@ Appendix A Some Identities of Polylogarithms
where $S(t)=\frac{e^{t/2}-e^{-t/2}}{t}$.
@@
$-\frac{1}{2}\operatorname{Res}\displaylimits_{q\to\frac{1}{\beta_{j}}}(y(q)-y(p))\,\frac{\sum_{\begin{subarray}{c}g_{1}+g_{2}=g\\ g_{i}>0\end{subarray}}\omega_{g_{1},1}^{\vee}(q)\,\omega_{g_{2},1}^{\vee}(q)}{dx(q)\,dy(q)}\,dx(q)$
@@
$\operatorname{Res}\displaylimits_{q\to\frac{1}{\beta_{j}}}(y(q)-y(p))\bigg{(}\int^{q}\omega_{g,1}^{\vee}-\frac{\frac{1}{2}\sum_{\begin{subarray}{c}g_{1}+g_{2}=g\\ g_{i}>0\end{subarray}}\omega_{g_{1},1}^{\vee}(q)\,\omega_{g_{2},1}^{\vee}(q)}{dx(q)\,dy(q)}\bigg{)}dx(q)$

### v2

@@ 3.1. Four punctured spheres
Depending on whether $|Q|<1$ or $|Q|>1$, one can write the sum in terms of $\mathrm{Li}_{3-2g}(Q)$ or $\mathrm{Li}_{3-2g}(Q^{-1})$ which are equal. However, as expansions in $Q$, only one of them is convergent, given the above choice. This was also observed in [78], giving two distinct chambers in complex structure moduli space. In this case, we infer  [^2] The bar on charges denote the anti D2-brane charges. The D-brane charges are classified by K-theory by the even integral cohomology of $\mathcal{X}$, $\Gamma\subset H^{\textrm{even}}(\mathcal{X},\mathbb{Z})$. A general charge vector is written as $\gamma\in\Gamma$, where $\gamma=(p^{0},p^{a},q_{a},q_{0})$, where focusing on a D2-branes that wraps a curve $C$ of class $\beta$ particularly $\gamma_{\textrm{D2}}=(0,0,\beta,\frac{1}{2}\beta\cdot c_{1}(\mathcal{X}))$. Since for CY3, $c_{1}(\mathcal{X})=0$, one has $\gamma_{D2}=-\gamma_{\overline{D2}}=(0,0,\beta,0)$. For all the examples we use this notation with overline for the anti-D2 branes.
@@
This matches with the results of [74, 78, 17].
@@
$Z_{\mathrm{DT}}=M(q)^{-2}\prod_{n=1}^{\infty}(1-Qq^{n})^{-n}$ (3.6)
equation (4.36) in [17].
@@ 3.2. Examples of five punctured spheres
Next, one could consider the case of gluing two copies of $\mathcal{O}(0,-2)$, corresponding to $s=2$ which corresponds to resolution of $\mathbb{C}\times\mathbb{C}^{2}/\mathbb{Z}_{3}$. In this case, as in [76], we consider $\beta_{1}=Q_{1},\beta_{2}=Q_{1}Q_{2}$. Looking at the genus $g$ free energy (2.15), one can check
@@
which matches the computations of [28, 68, 67, 66].
@@ 3.3. Examples of six punctured spheres
We take the example of [76] for the six-punctured sphere with $r=2,s=1$, with $\alpha_{1}=Q_{1},\alpha_{2}=Q_{1}Q_{2}Q_{3}$ and $\beta_{1}=Q_{1}Q_{2}$. The 5D indices can then be read off from (2.15) as
@@ Appendix A Some Identities of Polylogarithms
where $S(t)=\frac{e^{t/2}-e^{-t/2}}{t}$. We have also used $\big(z\frac{d}{dz}\big)^{2g}\log(1-az)=-\mathrm{Li}_{1-2g}(az)$, which follows directly from the definition of the polylogarithm.
@@
$-\frac{1}{2}\operatorname{Res}\displaylimits_{q\to\frac{1}{\beta_{j}}}(y(q)-y(p))\,\frac{\underset{\begin{subarray}{c}g_{1}+g_{2}=g,\\ g_{i}>0\end{subarray}}{\sum}\omega_{g_{1},1}^{\vee}(q)\,\omega_{g_{2},1}^{\vee}(q)}{dx(q)\,dy(q)}\,dx(q)$
@@
$\operatorname{Res}\displaylimits_{q\to\frac{1}{\beta_{j}}}(y(q)-y(p))\bigg(\int^{q}\omega_{g,1}^{\vee}-\frac{\frac{1}{2}\underset{\begin{subarray}{c}g_{1}+g_{2}=g,\\ g_{i}>0\end{subarray}}{\sum}\omega_{g_{1},1}^{\vee}(q)\,\omega_{g_{2},1}^{\vee}(q)}{dx(q)\,dy(q)}\bigg)dx(q)$
@@
conflict of interest statement
On behalf of all authors, the corresponding author states that there is no conflict of interest.